The density of materials changes when their temperature changes. the density of water is 1000kg/m cubed at 0 degrees Celsius, and is 700kg/m cubed at 300 degrees Celsius. calculate the volume of 5kg of water at 0 degrees Celsius.
1
Expert's answer
2018-05-23T11:56:08-0400
The volume of 5kg of water at 0 degrees Celsius is
V=m/ρ_0 =5/1000=0.005 m^3.